# Interacting Crystalline Topological Insulators in 2D with Spin-Orbit Coupling

### v1

@@ V Discussion
In summary, we have completed the classification of cFSPTs with TRS and U(1) charge-conservation for all the 17 wallpaper groups, and the result of the classification is summarized in Tab. 1. We have adopted a real-space based construction first proposed in Refs. 78, 79, 80, 81, 82, and we have introduced the concept of Wyckoff patches as a natural prescription to obtain the unit cell subdivision.
@@
Due to the constructive nature of the classification procedure we adopted, each term appearing in the products of Tab. 1 can be precisely associated to a topological index, and each index in turn characterizes the decorations of a specific Wyckoff patch. The indices corresponding to each entry of Tab. 1 are listed in App. B. For each wallpaper group the classification summarized in Tab. 1 always contains a $\mathbb{Z}$ factor counting the total charge in the unit cell, and a $\mathbb{Z}_{2}$ factor always appearing at the end (colored in magenta) for the intrinsic 2D topological phase, which does not require crystalline symmetries to be protected. In addition to those, the wallpaper cFSPTs are classified by a series of $\mathbb{Z}_{n}$ ($n=2,3,4,6$) terms (colored in black) that describe the charge distribution within the unit cell, and all the remaining $\mathbb{Z}_{2}$ factors (colored in cyan) classify intrinsically interacting decorations. The latter correspond to either mirror and rotation charges, or to the AIIM phase, namely the 1D non trivial cFSPT phase protected by mirror symmetry.
@@
We remark that in our classification scheme, two Hamiltonians are in the same cFSPT phase only if they can be deformed to one another without symmetry breaking or gap closing. In particular, we do not allow stacking with “ancilla” degrees of freedom. This is in contrast to the classification of the so-called strong topological phases where stacking with a product state does not change the topological phase. Depending on the choice of ancillas some of the entries in Table 1 can be trivialized. For instance, all $\mathbb{Z}$ entries corresponding to the U(1) charge per unit cell, can be changed by stacking with charged ancillas. These phases correspond to the so-called fragile topological phases [130, 131, 132, 133].
@@
The ability to associate indices to real-space decorations is a rather powerful aspect of the real-space construction, as it allows to not only extract the classification of cFSPTs, but also to construct fixed point wavefunctions that lie in each of the phases predicted by the classification. These correspond to wavefunctions built by gluing together decorations matching the values of the topological indices describing the phase. We have both recalled and proposed model Hamiltonians for the realization of all the decorations necessary towards the construction of cFSPTs as their unique ground state, and finally we discussed possible approaches towards distinguishing these phases.
@@
With this work, we hope to provide a tool that may support further study of interacting topological phases. In particular, we expect that this classification will be useful in characterizing more realistic Hamiltonian ground states that fall in the same symmetry classes as the cFSPTs analyzed here, namely the one of interacting spin-orbit coupled interacting insulators, that do not undergo spontaneous symmetry breaking. The model Hamiltonians presented here may be a starting point to test the resilience of cFSPTs against disorder and perturbation away from the fixed point wavefunctions. In addition, providing a classification to compare the results against, might be helpful in exploring the regime of validity of various methods and approaches used to detect and characterize cFSPTs.
@@ Acknowledgements.
We thank Apoorv Tiwari, Juven C. Wang, and Zheyan Wan for useful and extensive discussions in the early stages of this work. The authors acknowledge useful discussions with Max Geier and Christopher Mudry. M. O. S. acknowledges funding from the University of Zurich Forschungskredit (grant No. FK-23-110). Ö. M. A. is supported by Swiss National Science Foundation (SNSF) under Grant No. P500PT-214429 and National Science Foundation (NSF) DMR-2022428. T.N. and M. O. S. acknowledge support from the Swiss National Science Foundation (SNSF) through a Consolidator Grant (iTQC, TMCG-2-213805) and through Grant No. 200021E-198011 as part of the FOR 5249 (QUAST) lead by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).
@@ Appendix B Classification of wallpaper group crystalline SPTs
In this section we provide the missing steps followed to obtain the cFSPTs classification for all the 17 wallpaper groups, namely the various steps that leading to the entries in Tab. 1. Details on the wallpaper groups, such as the coordinates of the Wyckoff positions and their site-symmetry groups can be found for instance in Ref. 89, or online on the Bilbao Crystallographic Server website [134, 135, 136].
@@
Figure A5 depicts the unit cell and Wyckoff positions for each of the 17 wallpaper groups. In all the panels, each Wyckoff position is represented as a dot, line or shaded quadrilateral for 0D, 1D or 2D Wyckoff patches, respectively.
@@
Tables A1A17 list the classification of gappable decorations for the 17 wallpaper groups, prior to taking into account any equivalence relations. The caption of each table indicates the wallpaper group to which the Table refers to. For all of the Tabs. A1A17, the column labels indicate:
@@
    “Classification”: Classification for each Wyckoff patch, following the prescription of Tab. 3 in the main text, based on the entry in $G_{\text{spa}}$.
@@
    “Indices”: List of topological indices referred to the gappable decorations. Note that the order in which the indices are listed matches the order in which the groups are written in the product of groups in the column “Gappable decorations”.
For the topologial indices, we follow the convention introduced in the main text. We summarize the convention here for convenience. For a Wyckoff patch whose Wyckoff position is labeled by $wp$, we indicate by
@@
    $n_{wp}\in\mathbb{Z}$ (or $\mathbb{Z}_{n}$ with $n=2,3,4,6$ after the reduction due to equivalence relations): Number of Kramers’ pairs of 0D decorations at the Wyckoff patch $wp$;
@@
    $r_{wp}\in\mathbb{Z}^{\,}_{2}$: Rotation charge of 0D decorations at the 0D $wp$;
@@
    $m_{wp}\in\mathbb{Z}^{\,}_{2}$: Mirror charge of 0D decorations at the 0D or 1D Wyckoff patch $wp$;
@@
    $\nu_{wp}\in\mathbb{Z}^{\,}_{2}$: 1D topological phase of 1D decorations at the 1D Wyckoff patch $wp$. This index classifies whether the 1D decorations are in the trivial phase, or the intrinsically interacting nontrivial phase which we call AIIM phase, discussed in Sec. III.2 of the main text;
@@
When a list of such indices is written next to the product of groups the indices belong to, we keep the ordering consistent: For instance, the $i$th index listed in the column “Indices” belongs to the $i$th group of the product in “Gappable decorations”.
Note that while $n_{wp}$ and $x_{wp}$ are generically present for both interacting and non-interacting systems, instead the $r_{wp}$, $m_{wp}$, and $\nu_{wp}$ indices can assume nontrivial values only when interactions are present, i. e., their nontrivial values describe intrinsically interacting phases.
@@
In App. B.1, we list the equivalence relations for each of the 17 wallpaper groups, which allow to reduce the indices of the gappable decorations listed in Tabs A1A17. Equivalences are adiabatic and symmetry respecting deformations that allow to connect phases characterized by different topological indices without a gap closing occurring. These equivalences can either partially reduce, or fully trivialize the group to which the indices originally belonged.
@@
    Those obtained by symmetrically peeling off 0D decorations from a Wyckoff patch to another one, which only changes the value of the indices counting the number of Kramer’s pairs at the Wyckoff patches,
@@
After listing the equivalence processes, for every wallpaper group we then explicitly show the reduction of the indices occurring due to the equivalence processes. We write on the left hand side of the arrow $\rightarrow$” the original indices with their classification (omitting those indices that are left unchanged by all the equivalence relations), and on the right hand side the reduced classification. In particular, indices that are fully trivialized, namely whose group becomes $\mathbb{Z}_{1}$ after taking into account equivalence relations, are omitted on the right hand side of the arrow.
Table A18 lists the cFSPT classification obtained for each wallpaper group, alongside the indices corresponding to each factor in the classification. The first two columns identify the wallpaper group through its number (“No.”) and label (“Wallpaper”). The “Classification” column lists the full (interacting) classification, while the “Non-interacting Classification” column shows the classification retained in the absence of interactions. The last column, “Indices”, explicitly lists the topological indices that are associated to the (interacting) classification. Once more, the ordering of the indices listed in the “Indices” column and the groups appearing in the product in the “Classification” column is consistent.
Figure A5: Representation of the Wyckoff patches for the 17 wallpaper groups. The panels are labeled by the Wallpaper group’s number and label. Each Wyckoff patch is labeled by the Wyckoff position from which it originates, and it is represented according to its dimension: 0D Wyckoff patches are marked by dots, 1D by dashed or dotted lines, and 2D by shaded quadrilateral regions. In each panel, a single unit cell is marked by black lines. For the wallpaper groups No. 14, 15 and 17, dashed and dotted lines showcase how some of the 1D patches can be decorated in different ways: 1D decorations can be glued along dashed lines only, dotted lines only, or along both at the same time.
@@ B.1 Enumeration of gappable decorations and equivalences
No. Wallpaper Classification Non-interacting Classification Indices
@@
Table A18: Classification of cFSPTs in wallpaper groups. Every wallpaper group has a $\mathbb{Z}$ factor in its classification due to the total number of Kramers’ pairs in the unit cell (with index $n$), and a $\mathbb{Z}_{2}$ factor (with index $x_{wp}$, with $wp$ the general Wyckoff position label) deriving from the intrinsic 2D SPT phase that does not require crystalline symmetries to be realized (colored in magenta). Terms in the classification and indices colored in black correspond to phases that exist both in the absence or presence of interactions. The terms corresponding to intrinsically interacting phases in the classification and their corresponding indices are highlighted in cyan. Note that the “Indices” column refers to the “Classification” column, and the ordering is consistent between the two.

### v2

@@ V Discussion
In summary, we have completed the classification of cFSPTs with TRS and U(1) charge-conservation for all the 17 wallpaper groups, and the result of the classification is summarized in Table 1. We have adopted a real-space based construction first proposed in Refs. 78, 79, 80, 81, 82, and we have introduced the concept of Wyckoff patches as a natural prescription to obtain the unit-cell subdivision.
@@
Due to the constructive nature of the classification procedure we adopted, each term appearing in the products of Table 1 can be precisely associated to a topological index, and each index in turn characterizes the decorations of a specific Wyckoff patch. The indices corresponding to each entry of Table 1 are listed in Appendix B. For each wallpaper group the classification summarized in Table 1 always contains a $\mathbb{Z}$ factor counting the total charge in the unit-cell, and a $\mathbb{Z}_{2}$ factor always appearing at the end (colored in magenta) for the intrinsic 2D topological phase, which does not require crystalline symmetries to be protected. In addition to those, the wallpaper cFSPTs are classified by a series of $\mathbb{Z}_{n}$ ($n=2,3,4,6$) terms (colored in black) that describe the charge distribution within the unit-cell, and all the remaining $\mathbb{Z}_{2}$ factors (colored in cyan) classify intrinsically interacting decorations. The latter correspond to either mirror and rotation charges, or to the AIIM phase, namely, the 1D non trivial cFSPT phase protected by mirror symmetry.
@@
We remark that in our classification scheme, two Hamiltonians are in the same cFSPT phase only if they can be deformed to one another without symmetry breaking or gap closing. In particular, we do not allow stacking with “ancilla” degrees of freedom. This is in contrast to the classification of the so-called strong topological phases where stacking with a product state does not change the topological phase. Depending on the choice of ancillas some of the entries in Table 1 can be trivialized. For instance, all $\mathbb{Z}$ entries corresponding to the U(1) charge per unit-cell, can be changed by stacking with charged ancillas. These phases correspond to the so-called fragile topological phases [131, 132, 133, 134].
@@
The ability to associate indices to real-space decorations is a rather powerful aspect of the real-space construction, as it allows to not only extract the classification of cFSPTs, but also to construct fixed-point wave functions that lie in each of the phases predicted by the classification. These correspond to wave functions built by gluing together decorations matching the values of the topological indices describing the phase. We have both recalled and proposed model Hamiltonians for the realization of all the decorations necessary towards the construction of cFSPTs as their unique ground state, and finally we discussed possible approaches towards distinguishing these phases.
@@
With this work, we hope to provide a tool that may support further study of interacting topological phases. In particular, we expect that this classification will be useful in characterizing more realistic Hamiltonian ground states that fall in the same symmetry classes as the cFSPTs analyzed here, namely, the one of interacting spin-orbit coupled interacting insulators, that do not undergo spontaneous symmetry breaking. The model Hamiltonians presented here may be a starting point to test the resilience of cFSPTs against disorder and perturbation away from the fixed-point wave functions. In addition, providing a classification to compare the results against, might be helpful in exploring the regime of validity of various methods and approaches used to detect and characterize cFSPTs.
@@ Acknowledgements.
We thank Apoorv Tiwari, Juven C. Wang, and Zheyan Wan for useful and extensive discussions in the early stages of this work. The authors acknowledge useful discussions with Max Geier and Christopher Mudry. M. O. S. acknowledges funding from the University of Zurich Forschungskredit (Grant No. FK-23-110). Ö. M. A. is supported by Swiss National Science Foundation (SNSF) under Grant No. P500PT-214429 and National Science Foundation (NSF) Grant No. DMR-2022428. T.N. and M. O. S. acknowledge support from the Swiss National Science Foundation (SNSF) through a Consolidator Grant (iTQC, TMCG-2-213805) and through Grant No. 200021E-198011 as part of the FOR 5249 (QUAST) lead by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).
Data availability
This publication is theoretical work that does not require supporting research data.
@@ Appendix B Classification of wallpaper group crystalline SPTs
In this appendix we provide the missing steps followed to obtain the cFSPTs classification for all the 17 wallpaper groups, namely, the various steps that lead to the entries in Table 1. Details on the wallpaper groups, such as the coordinates of the Wyckoff positions and their site-symmetry groups, can be found for instance in Ref. 89, or online on the Bilbao Crystallographic Server website [135, 136, 137].
@@
Figure A5 depicts the unit-cell and Wyckoff positions for each of the 17 wallpaper groups. In all the panels, each Wyckoff position is represented as a dot, line or shaded quadrilateral for 0D, 1D or 2D Wyckoff patches, respectively.
@@
Tables A1A17 list the classification of gappable decorations for the 17 wallpaper groups, prior to taking into account any equivalence relations. The caption of each table indicates the wallpaper group to which the table refers to. For all of the Tables  A1A17, the column labels indicate the following:
@@
    “Classification”: Classification for each Wyckoff patch, following the prescription of Table 3 in the main text, based on the entry in $G_{\text{spa}}$.
@@
    “Indices”: List of topological indices referred to the gappable decorations. Note that the order in which the indices are listed matches the order in which the groups are written in the product of groups in the column “Gappable decorations.”
For the topological indices, we follow the convention introduced in the main text. We summarize the convention here for convenience. For a Wyckoff patch whose Wyckoff position is labeled by $wp$, we indicate by the follwing:
@@
    $n_{wp}\in\mathbb{Z}$ (or $\mathbb{Z}_{n}$ with $n=2,3,4,6$ after the reduction due to equivalence relations): Number of Kramers’ pairs of 0D decorations at the Wyckoff patch $wp$.
@@
    $r_{wp}\in\mathbb{Z}^{\,}_{2}$: Rotation charge of 0D decorations at the 0D $wp$.
@@
    $m_{wp}\in\mathbb{Z}^{\,}_{2}$: Mirror charge of 0D decorations at the 0D or 1D Wyckoff patch $wp$.
@@
    $\nu_{wp}\in\mathbb{Z}^{\,}_{2}$: 1D topological phase of 1D decorations at the 1D Wyckoff patch $wp$. This index classifies whether the 1D decorations are in the trivial phase, or the intrinsically interacting nontrivial phase which we call AIIM phase, discussed in Sec. III.2 of the main text.
@@
When a list of such indices is written next to the product of groups the indices belong to, we keep the ordering consistent: For instance, the $i$th index listed in the column “Indices” belongs to the $i$th group of the product in “Gappable decorations.”
Note that while $n_{wp}$ and $x_{wp}$ are generically present for both interacting and noninteracting systems, instead the $r_{wp}$, $m_{wp}$, and $\nu_{wp}$ indices can assume nontrivial values only when interactions are present, i. e., their nontrivial values describe intrinsically interacting phases.
@@
In Appendix B.1, we list the equivalence relations for each of the 17 wallpaper groups, which allow to reduce the indices of the gappable decorations listed in Tables A1A17. Equivalences are adiabatic and symmetry respecting deformations that allow to connect phases characterized by different topological indices without a gap closing occurring. These equivalences can either partially reduce, or fully trivialize, the group to which the indices originally belonged.
@@
    Those obtained by symmetrically peeling off 0D decorations from a Wyckoff patch to another one, which only changes the value of the indices counting the number of Kramer’s pairs at the Wyckoff patches.
@@
After listing the equivalence processes, for every wallpaper group we then explicitly show the reduction of the indices occurring due to the equivalence processes. We write on the left-hand side of the arrow $\rightarrow$” the original indices with their classification (omitting those indices that are left unchanged by all the equivalence relations), and on the right hand side the reduced classification. In particular, indices that are fully trivialized, namely, whose group becomes $\mathbb{Z}_{1}$ after taking into account equivalence relations, are omitted on the right hand side of the arrow.
Classification of cFSPTs in wallpaper groups with corresponding indices.
Table A18 lists the cFSPT classification obtained for each wallpaper group, alongside the indices corresponding to each factor in the classification. The first two columns identify the wallpaper group through its number (“No.”) and label (“Wallpaper”). The “Classification” column lists the full (interacting) classification, while the “Noninteracting classification” column shows the classification retained in the absence of interactions. The last column, “Indices”, explicitly lists the topological indices that are associated to the (interacting) classification. Once more, the ordering of the indices listed in the “Indices” column and the groups appearing in the product in the “Classification” column is consistent.
Figure A5: Representation of the Wyckoff patches for the 17 wallpaper groups. The panels are labeled by the Wallpaper group’s number and label. Each Wyckoff patch is labeled by the Wyckoff position from which it originates, and it is represented according to its dimension: 0D Wyckoff patches are marked by dots, 1D by dashed or dotted lines, and 2D by shaded quadrilateral regions. In each panel, a single unit-cell is marked by black lines. For the wallpaper groups Nos. 14, 15 and 17, dashed and dotted lines showcase how some of the 1D patches can be decorated in different ways: 1D decorations can be glued along dashed lines only, dotted lines only, or along both at the same time.
@@ B.1 Enumeration of gappable decorations and equivalences
In this section, we list all the equivalence relations for each 993 of the wallpaper groups [Sec. B].
@@
Finally, Table A18 shows the classification for cFSPTs, for each of the 17 wallpaper groups in two dimensions.
No. Wallpaper Classification Noninteracting classification Indices
@@
Table A18: Classification of cFSPTs in wallpaper groups. Every wallpaper group has a $\mathbb{Z}$ factor in its classification due to the total number of Kramers’ pairs in the unit-cell (with index $n$), and a $\mathbb{Z}_{2}$ factor (with index $x_{wp}$, with $wp$ the general Wyckoff position label) deriving from the intrinsic 2D SPT phase that does not require crystalline symmetries to be realized (colored in magenta). Terms in the classification and indices colored in black correspond to phases that exist both in the absence or presence of interactions. The terms corresponding to intrinsically interacting phases in the classification and their corresponding indices are highlighted in cyan. Note that the “Indices” column refers to the “Classification” column, and the ordering is consistent between the two.